OpenStudy (anonymous):

question below:

OpenStudy (anonymous):

using the substitution \[y = 2e ^{x}\] solve: \[e ^{x} = 7 - 12e ^{-x}\]

hartnn (hartnn):

multiply both sides by e^x u get \(e^{(2x)}=7e^x-12 \\ now \quad e^x=y/2 \\ e^{2x}=y^2/4 \\ y^2/4-7(y/2)+12=0\) can u solve this quadratic in y ??

hartnn (hartnn):

u get \(y^2-14x+48=0\) which is easily factorable.

OpenStudy (anonymous):

hold on.. in the first step.. why would you multiply both sides by e^x?

OpenStudy (anonymous):

I am SO sorry. Typo! the substitution had to be y= e^x. NOT 2e^x

hartnn (hartnn):

i would multiply both sides by e^x to get the e^{2x} term for substitution \(e^{(2x)}=7e^x-12 \\ now \quad e^x=y \\ e^{2x}=y^2 \\ y^2-7(y)+12=0\)

hartnn (hartnn):

still easily factorable........

OpenStudy (anonymous):

Okay yes! I get it! Thank you so much for your help! :)